OpenStudy (anonymous):

4(y – 4) = 8 A) –2 B) 2 C) 4 D) 6 Thanks :D

OpenStudy (anonymous):



OpenStudy (mayankdevnani):

4y-16=8 4y=8+16 4y=24 y=6

OpenStudy (anonymous):

well, first u multiply across the parenthesis...4y-16=8 now u hav an equation, so u add 16 and then divide by 4...and that will give u an answer :)

OpenStudy (anonymous):

\[4(y-4)=8\]\[y-4=\frac{ 8 }{ 4 }=2\]\[y-4+4=2+4\]\[y=6\]

OpenStudy (mayankdevnani):

ok @Hailey!